# Physics-informed machine learning for combustion: A review

### v2

@@ 5.3 Turbulent combustion and experimental data
Compared to laminar combustion and using simulation data for learning or validating, the problems are more complex for turbulent combustion and when only practical experimental data are available due to high uncertainty of these situations. As an early attempt, Nandi et al. 327, 328 used PINNs to develop digital twins for 2D and 3D industry-scale boiler combustion without observation data (forward problems). The zero-equation model was used for the turbulent Reynolds stresses 327 and various reaction source terms were tested, including the linear form of two species 327 and one-step mechanism of CH4 328. Some strategies were proposed to handle the coupling between the source terms of $T$ and $Y$. Von Saldern et al. 229 applied PINNs to the assimilation of turbulent mean flow data of axisymmetric jet flows. In their first example, the mean density fields of a turbulent jet flame were reconstructed based on the data of velocity fields, which were generated by LES. Although only the continuity equation was considered in the physical loss, the velocity and density fields were accurately reconstructed. Peng et al. 311 used PINNs to reconstruct the heat flow field of Sandia Flame D 329 based on the data of velocity and temperature, which were obtained by simulation with the partially premixed combustion model. The predicted variables and physical losses did not consider the species and chemical reactions, but considered the turbulence kinetic energy ($k$), its dissipation rate ($\varepsilon$), and their transport equations, which is a common turbulence model. The results showed that PINNs could accurately reconstruct the fields within both small and large time intervals. Compared to data-driven methods, PINNs was shown to yield more accurate results and require less data when accuracy requirements are the same.
In addition to mean and filtered turbulence, PINNs can also be applied to direct turbulence, directly integrating the original reacting flow equations, i.e., Equations (3.38) to (3.42). Liu et al. 310 used PINNs to reconstruct the fields of 3D turbulent flames based on sparse, noisy data sampled from high-fidelity DNS data. Three cases were tested: freely-propagating planar premixed flames with low and high levels of turbulence, and slot-jet flames with a jet Reynolds number of 850. For the physical models, temperature-dependent properties, constant Prandtl number, and unity Lewis number were considered. A two-step mechanism of CH4 was adopted, which consisted of six species and two reactions. Three PINN sub-models were employed to learn momentum, species, and temperature equations, respectively, and the species conservation was also considered in the losses (Equation (3.1)). Self-adaptive linear layers were utilized and a warmup pretraining was conducted to facilitate model convergence. It is shown that the model can accurately reconstruct the fields and get satisfactory results of flow modes, vortical structures, and PDFs of variables and flame curvature, and the results were more accurate than the linear interpolation method, as shown in Figure 20. Subsequently, Liu et al. 304 applied NNs to 3D reconstruction of temperature and velocity from 2D measurements of turbulent flames. First, the 3D temperature field were reconstructed from 2D data using data-driven models: two vector-quantized VAEs first learned the latent spaces of 2D and 3D temperature fields separately, and then a DM parameterized by an unconstrained transformer learned the 2D to 3D translation using the learned vector-quantized space. After that, a wavenumber-based PINN (WN-PINN) was proposed to reconstruct the 3D velocity fields based on the 2D velocity and reconstructed 3D temperature data, where momentum equation was used as the physical constraint. To address the multiscale issue, the velocity field was decomposed into low and high wavenumber components and learned by two separate NNs. Two cases were tested: freely propagating planar premixed flames with three operating conditions, and a swirling premixed flame. Results showed excellent reconstruction for both temperature and velocity. More importantly, the results showed that for unseen cases, both pre-trained models and models trained on these new cases from scratch are not as accurate as the fine-tuned pre-trained models, demonstrating the effectiveness and efficiency of transfer learning.
@@
Figure 20: PINNs reconstruct turbulent flames: comparisons of the instantaneous 3D distributions of $Y_{\mathrm{O}_{2}}$ from linear interpolation, PINNs and DNS data. Reprinted from Reference 310, with permission.
@@
For experimental combustion, there are discrepancies between the data and pre-specified physical models, especially for turbulence and reaction models. As shown in Table 5, two strategies are adopted in existing studies to deal with this issue: (1) Reduce or simplify the models; (2) Learn the models, i.e., closure terms. Taassob et al. 312 used PINNs to evaluate closure terms for turbulence and chemical source terms in the Sandia turbulent nonpremixed flames 329. As illustrated in Figure 21, the NN took coordinates ($x$, $r$) and Reynolds number as input for parametric learning, and output not only the mean flow fields, including $\overline{\rho}$, $\widetilde{\mathbf{U}}$, $\overline{p}$, $\widetilde{Z}$, and two principal components ($\widetilde{\mathrm{PC}}$) of $T$ and $Y$, but also the Reynolds stress ($\widetilde{u^{\prime}v^{\prime}}$) and turbulent viscosity ($\mu_{t}$) to learn the turbulence closure, which was achieved by adding the $\mu_{t}$ relation to the physical losses:
@@
Figure 21: Schematic of the PINNs for turbulent combustion closure. Reprinted from Reference 312, with permission.
Liu et al. 110 used PINNs to reconstruct the soot fields in acoustically forced laminar sooting flames in two steps. In the first step, since the velocity data measured by particle image velocimetry (PIV) were missing in sooting regions, and the temperature data measured by non-linear excitation regime two-line atomic fluorescence (nTLAF) were missing in low-temperature regions, a PINN was used to reconstruct the velocity and temperature fields. The momentum and temperature equations were used as physical losses where the heat release and radiation were neglected due to low temperatures. Results showed that PINN could learn the available data well and reconstruct the missing fields. Next, a physics-informed AE-based DeepONet (PI-AE-DeepONet) with two AEs was used to reconstruct the soot fields, as shown in Figure 22. The simplified transport equation of soot mass fraction was taken as the physical loss where the unknown source term was learned by the second AE. The first AE was used to learn the latent features of OH images and PINN-reconstructed $T$ images via supervised learning. The DeepONet took the latent features as the branch inputs and output the soot volume fraction. It was shown that PI-AE-DeepONet yielded better predictions than various data-driven models, including U-Net, FNO, the first AE, and AE-DeepONet, and could learn the source terms accurately.
Figure 22: Schematic of the AE-based DeepONet framework to reconstruct the soot fields. Reprinted from Reference 110, with permission.
Yadav et al. 313 developed the reactive flow PINNs (RF-PINNs) to reconstruct all flow quantities of interest solely based on sparse velocity or temperature data. The framework of RF-PINNs is given in Figure 23. Three cases were tested: 1D freely-propagating flames, 2D laminar slit flames, and 2D experimental turbulent flames. Based on some assumptions, the equations of $T$ and $Y_{k}$ were reduced to the equation of the progress variable $c$, which was denoted as $\theta$ in the original paper. For laminar cases, the reaction source term of $c$ was pre-learned by a data-driven NN: $\dot{\omega}_{c}=\mathrm{NN}(c)$, while for turbulent cases, the source term was computed by the eddy break-up (EBU) model:
@@
Figure 23: The framework of RF-PINNs. Reprinted from Reference 313.
Discrete PINNs have also been applied to turbulent combustion. Sharma et al. 119 applied the PIML approach to predict the dynamic behaviour of hydrogen jet flames in an AR manner (see Section 2.3.1). As illustrated in Figure 24, the physics-informed hybrid multiscale and partitioned network (PiMAPNet) was constructed that integrated low-resolution physics-based models (numerical solver) with trainable NNs, so it can be viewed as a hybrid solver. The coarse grid numerical solver served as the physics-informed component of the model, which was similar to the physics-informed way in References 117, 118, 116 that integrated a numerical solver. Three other methods were selected for comparison: a coarse grid numerical solver without ML component, a purely data-driven CNN without physics-informed component, and a PPNN variant. As mentioned in Section 2.3.4, the original PPNN preserves/embeds the PDEs by discretizing the PDE RHS using convolutions filters instead of a numerical solver. Compared to the PPNN variant, PiMAPNet decomposed the state vector into hydrodynamic/flow variables and thermochemical variables. The flow variables were learned by a single CNN while the thermochemical variables were learned by a mixture-of-experts (MoE) that partitioned the thermochemical state-space so that separate CNNs could be assigned to different regions of the manifold. Results showed that the data-driven CNN yielded unphysical values and the PPNN variant exhibited the hotspot issue. In contrast, PiMAPNet successfully mitigated the issues of bias accumulation and hotspots and showed superior extrapolation performance. It was shown that the MoE effectively partitioned the manifold in a physically meaningful way, distinctly isolating the fuel-rich, fuel-lean, and mixing regions, which was the core reason for its superior performance. Speed analysis also showed that PiMAPNet was faster than a numerical solver with comparable accuracies.
Figure 24: The architecture of PiMAPNet. Reprinted from Reference 119, with permission.
In addition to AR learning, discrete PINNs can also be applied to field reconstruction. Based on the multi-receptive-field CNN 302, Zhang et al. 303 developed the combustion reacting flow PINNs (CRF-PINNs) for field reconstruction of laminar and turbulent combustion. As shown in Figure 25, CRF-PINNs can be used in three typical scenarios of field reconstruction: cross-field translation, SR reconstruction, and denoising, where the reacting flow equations in a cylindrical coordinate system and the species conservation serve as the physical constraints, eliminating the need for high-fidelity labeled data during training. For the laminar Bunsen flame, with one of the velocity, temperature, and species fields, CRF-PINNs can accurately reconstruct the other two. The cross-field translation was also demonstrated for unsteady Bunsen flames. For SR reconstruction and denoising, CRF-PINNs showed superior performance over the corresponding data-driven baseline models. For turbulent flames, the turbulent viscosity and the turbulence-chemistry interaction coefficient are learned by the NN along with unknown high-fidelity fields. The Sandia flame F 329 was tested for sparse reconstruction from partial, experimental measurements. The 13 measurement points along the centerline were used for evaluation, demonstrating satisfactory model accuracy. In summary, this work proves the feasibility of discrete PINNs for various scenarios of field reconstruction, and again showcases the treatment of turbulence and chemistry in PINNs.
Figure 25: CRF-PINN for three typical scenarios of field reconstruction in combustion reacting flows: (a) cross-field translation, (b) SR reconstruction, and (c) denoising. Reprinted from Reference 303, with permission.
@@ 5.4 Supersonic combustion
For supersonic combustion, the core feature is the presence of shock waves and their interaction with combustion, which poses challenges for experimental implementation, numerical solutions, and PINNs. In the early work of Mao et al. 330, they applied the DeepM&Mnet 331 for hypersonics to predict the coupled flow and finite-rate chemistry behind a normal shock. DeepM&Mnet was proposed as a data assimilation framework for fast simulating multiphysics and multiscale problems, where the pre-trained DeepONets were used as building blocks to constrain the solutions 331. In Reference 330, the global mass conservation was constrained by adding a loss term, thus making the framework physics-informed. Comparative analysis demonstrated the effectiveness of this constraint on accuracy improvement. Guo et al. 332 developed the physics-informed residual CNNs to reconstruct the flow field at the inlet of scramjet with the Mach number $Ma=10$, where the flow was non-viscous and non-reacting. After this study, Deng et al. 315 developed a physics-informed U-shaped residual NN (PI-ResUNet) to reconstruct the unsteady combustion flow field in a scramjet combustion chamber with $Ma=2$. As shown in Figure 26, the 2D unsteady compressible NS equations without reaction source terms were taken as the physical constraints and the operating parameters (Mach number and effective viscosity coefficient) were taken as the inputs to form a parametric learning. Comparing the PI-ResUNet, ResNet, PI-ResNet, and PI-UNet, it was demonstrated that introducing the physical constrains and residual structures could effectively improve the model robustness. Note that although the hidden layers of PI-ResUNet are convolutional layers, the inputs and outputs are still linear layers. Therefore, it is still a continuous PINN that uses AD for differentiation instead of the convolutional differentiation adopted in discrete PINNs (Section 2.3.1).
@@
Figure 26: The framework of PI-ResUNet for reconstructing the combustion flow field in a scramjet combustor, where $r$ indicates the density ($\rho$) and $\mu$ includes both the turbulent and laminar viscosity coefficients. Reprinted from Reference 315, with permission.
@@
Wang et al. 314 applied PINNs to reconstruct the combustion field in rotating detonation combustors (RDCs) based on partially observed data. As shown in Figure 27, a PINN was built to describe the 1D RDC numerical model and estimate the continuous physical fields, taking the space-time coordinates as inputs and physical fields as outputs. In this study, the observations included only ten sets of points with fixed spatial coordinates, representing measured data with fixed sensors. The studied problem was an inverse problem, where the equations and partial true results were known, but the overall solution of the equations was unknown. Results showed that compared to interpolation-based methods, PINN could better capture different distribution patterns for the flow characteristics before and after the detonation wave, which matched the ground truth satisfactorily. The proposed PINN can serve as a surrogate model of the RDC, which has promising engineering value in the design optimization and real-time monitoring of RDCs. After this study, Wang et al. 318, 319 further used the discrete PINNs for SR in RDCs, which will be discussed in Section 5.5.
Figure 27: Illustration of the flow field reconstruction problem for RDCs. Reprinted from Reference 314, with permission.

### v3

@@ 5.3 Turbulent combustion and experimental data
Compared to laminar combustion and using simulation data for learning or validating, the problems are more complex for turbulent combustion and when only practical experimental data are available due to high uncertainty of these situations. Some studies on PINNs for turbulent combustion and experimental data are shown in Figure 9.
@@
Figure 9: PINNs for turbulent combustion and experimental data. (a) PINNs reconstruct turbulent flames: comparisons of the instantaneous 3D distributions of $Y_{\mathrm{O}_{2}}$ from linear interpolation, PINNs and DNS data 307. (b) PINNs for turbulent combustion closure 311. (c) The AE-based DeepONet framework to reconstruct the soot fields 110. (d) The framework of RF-PINNs 312. (e) The architecture of PiMAPNet 119. Figures are reprinted with permission.
The most straightforward way to apply PINNs to turbulent combustion is directly integrating the original reacting flow equations, i.e., Equations (3.38) to (3.42), without any turbulence models or turbulent combustion models. Liu et al. 307 used PINNs to reconstruct the fields of 3D turbulent flames based on sparse, noisy data sampled from high-fidelity DNS data. Three cases were tested: freely-propagating planar premixed flames with low and high levels of turbulence, and slot-jet flames with a jet Reynolds number of 850. For the physical models, temperature-dependent properties, constant Prandtl number, and unity Lewis number were considered. A two-step mechanism of CH4 was adopted, which consisted of six species and two reactions. Three PINN sub-models were employed to learn momentum, species, and temperature equations, respectively, and the species conservation was also considered in the losses (Equation (3.1)). Self-adaptive linear layers were utilized and a warmup pretraining was conducted to facilitate model convergence. It is shown that the model can accurately reconstruct the fields and get satisfactory results of flow modes, vortical structures, and PDFs of variables and flame curvature, and the results were more accurate than the linear interpolation method, as shown in Figure 9(a). Subsequently, Liu et al. 308 applied NNs to 3D reconstruction of temperature and velocity from 2D measurements of turbulent flames. First, the 3D temperature field were reconstructed from 2D data using data-driven models: two vector-quantized VAEs first learned the latent spaces of 2D and 3D temperature fields separately, and then a DM parameterized by an unconstrained transformer learned the 2D to 3D translation using the learned vector-quantized space. After that, a wavenumber-based PINN (WN-PINN) was proposed to reconstruct the 3D velocity fields based on the 2D velocity and reconstructed 3D temperature data, where momentum equation was used as the physical constraint. To address the multiscale issue, the velocity field was decomposed into low and high wavenumber components and learned by two separate NNs. Two cases were tested: freely propagating planar premixed flames with three operating conditions, and a swirling premixed flame. Results showed excellent reconstruction for both temperature and velocity. More importantly, the results showed that for unseen cases, both pre-trained models and models trained on these new cases from scratch are not as accurate as the fine-tuned pre-trained models, demonstrating the effectiveness and efficiency of transfer learning.
In addition to direct turbulence, PINNs have also been widely applied to turbulent combustion assuming mean and filtered turbulence. As an early attempt, Nandi et al. 327, 328 used PINNs to develop digital twins for 2D and 3D industry-scale boiler combustion without observation data (forward problems). The zero-equation model was used for the turbulent Reynolds stresses 327 and various reaction source terms were tested, including the linear form of two species 327 and one-step mechanism of CH4 328. Some strategies were proposed to handle the coupling between the source terms of $T$ and $Y$. Von Saldern et al. 229 applied PINNs to the assimilation of turbulent mean flow data of axisymmetric jet flows. In their first example, the mean density fields of a turbulent jet flame were reconstructed based on the data of velocity fields, which were generated by LES. Although only the continuity equation was considered in the physical loss, the velocity and density fields were accurately reconstructed. Peng et al. 309 used PINNs to reconstruct the heat flow field of Sandia Flame D 329 based on the data of velocity and temperature, which were obtained by simulation with the partially premixed combustion model. The predicted variables and physical losses did not consider the species and chemical reactions, but considered the turbulence kinetic energy ($k$), its dissipation rate ($\varepsilon$), and their transport equations, which is a common turbulence model. The results showed that PINNs could accurately reconstruct the fields within both small and large time intervals. Compared to data-driven methods, PINNs was shown to yield more accurate results and require less data when accuracy requirements are the same. Cao et al. 310 applied PINNs for modeling 2D steady turbulent combustion of CH4 and air, where the $k-\varepsilon$ model and eddy dissipation model (EDM) were adopted for modeling turbulence and reaction rates, respectively. For forward problems, PINNs could successfully simulate the non-reacting jet flows but exhibited noticeable discrepancies from reference fields for reacting cases, demonstrating the significant challenges of forwardly solving turbulent combustion problems using PINNs. With a small amount of velocity data, the PINN results achieved a high degree of accuracy. The superiority of PINNs over traditional methods for surrogate modeling was also validated by the results of parametric modeling.
@@
For experimental combustion, there are discrepancies between the data and pre-specified physical models, especially for turbulence and reaction models. As shown in Table 5, two strategies are adopted in existing studies to deal with this issue: (1) Reduce or simplify the models; (2) Learn the models, i.e., closure terms. Taassob et al. 311 used PINNs to evaluate closure terms for turbulence and chemical source terms in the Sandia turbulent nonpremixed flames 329. As illustrated in Figure 9(b), the NN took coordinates ($x$, $r$) and Reynolds number as input for parametric learning, and output not only the mean flow fields, including $\overline{\rho}$, $\widetilde{\mathbf{U}}$, $\overline{p}$, $\widetilde{Z}$, and two principal components ($\widetilde{\mathrm{PC}}$) of $T$ and $Y$, but also the Reynolds stress ($\widetilde{u^{\prime}v^{\prime}}$) and turbulent viscosity ($\mu_{t}$) to learn the turbulence closure, which was achieved by adding the $\mu_{t}$ relation to the physical losses:
@@
Liu et al. 110 used PINNs to reconstruct the soot fields in acoustically forced laminar sooting flames in two steps. In the first step, since the velocity data measured by particle image velocimetry (PIV) were missing in sooting regions, and the temperature data measured by non-linear excitation regime two-line atomic fluorescence (nTLAF) were missing in low-temperature regions, a PINN was used to reconstruct the velocity and temperature fields. The momentum and temperature equations were used as physical losses where the heat release and radiation were neglected due to low temperatures. Results showed that PINN could learn the available data well and reconstruct the missing fields. Next, a physics-informed AE-based DeepONet (PI-AE-DeepONet) with two AEs was used to reconstruct the soot fields, as shown in Figure 9(c). The simplified transport equation of soot mass fraction was taken as the physical loss where the unknown source term was learned by the second AE. The first AE was used to learn the latent features of OH images and PINN-reconstructed $T$ images via supervised learning. The DeepONet took the latent features as the branch inputs and output the soot volume fraction. It was shown that PI-AE-DeepONet yielded better predictions than various data-driven models, including U-Net, FNO, the first AE, and AE-DeepONet, and could learn the source terms accurately.
Yadav et al. 312 developed the reactive flow PINNs (RF-PINNs) to reconstruct all flow quantities of interest solely based on sparse velocity or temperature data. The framework of RF-PINNs is given in Figure 9(d). Three cases were tested: 1D freely-propagating flames, 2D laminar slit flames, and 2D experimental turbulent flames. Based on some assumptions, the equations of $T$ and $Y_{k}$ were reduced to the equation of the progress variable $c$, which was denoted as $\theta$ in the original paper. For laminar cases, the reaction source term of $c$ was pre-learned by a data-driven NN: $\dot{\omega}_{c}=\mathrm{NN}(c)$, while for turbulent cases, the source term was computed by the eddy break-up (EBU) model:
@@
Discrete PINNs have also been applied to turbulent combustion, including field reconstruction and AR learning. For field reconstruction, especially for SR reconstruction, details will be introduced in Section 5.5. For AR learning, Sharma et al. 119 applied the PIML approach to predict the dynamic behaviour of hydrogen jet flames in an AR manner (see Section 2.3.1). As illustrated in Figure 9(e), the physics-informed hybrid multiscale and partitioned network (PiMAPNet) was constructed that integrated low-resolution physics-based models (numerical solver) with trainable NNs, so it can be viewed as a hybrid solver. The coarse grid numerical solver served as the physics-informed component of the model, which was similar to the physics-informed way in References 117, 118, 116 that integrated a numerical solver. Three other methods were selected for comparison: a coarse grid numerical solver without ML component, a purely data-driven CNN without physics-informed component, and a PPNN variant. As mentioned in Section 2.3.4, the original PPNN preserves/embeds the PDEs by discretizing the PDE RHS using convolutions filters instead of a numerical solver. Compared to the PPNN variant, PiMAPNet decomposed the state vector into hydrodynamic/flow variables and thermochemical variables. The flow variables were learned by a single CNN while the thermochemical variables were learned by a mixture-of-experts (MoE) that partitioned the thermochemical state-space so that separate CNNs could be assigned to different regions of the manifold. Results showed that the data-driven CNN yielded unphysical values and the PPNN variant exhibited the hotspot issue. In contrast, PiMAPNet successfully mitigated the issues of bias accumulation and hotspots and showed superior extrapolation performance. It was shown that the MoE effectively partitioned the manifold in a physically meaningful way, distinctly isolating the fuel-rich, fuel-lean, and mixing regions, which was the core reason for its superior performance. Speed analysis also showed that PiMAPNet was faster than a numerical solver with comparable accuracies.
@@ 5.4 Supersonic combustion
For supersonic combustion, the core feature is the presence of shock waves and their interaction with combustion, which poses challenges for experimental implementation, numerical solutions, and PINNs. In the early work of Mao et al. 330, they applied the DeepM&Mnet 331 for hypersonics to predict the coupled flow and finite-rate chemistry behind a normal shock. DeepM&Mnet was proposed as a data assimilation framework for fast simulating multiphysics and multiscale problems, where the pre-trained DeepONets were used as building blocks to constrain the solutions 331. In Reference 330, the global mass conservation was constrained by adding a loss term, thus making the framework physics-informed. Comparative analysis demonstrated the effectiveness of this constraint on accuracy improvement. Guo et al. 332 developed the physics-informed residual CNNs to reconstruct the flow field at the inlet of scramjet with the Mach number $Ma=10$, where the flow was non-viscous and non-reacting. After this study, Deng et al. 314 developed a physics-informed U-shaped residual NN (PI-ResUNet) to reconstruct the unsteady combustion flow field in a scramjet combustion chamber with $Ma=2$. As shown in Figure 10(a), the 2D unsteady compressible NS equations without reaction source terms were taken as the physical constraints and the operating parameters (Mach number and effective viscosity coefficient) were taken as the inputs to form a parametric learning. Comparing the PI-ResUNet, ResNet, PI-ResNet, and PI-UNet, it was demonstrated that introducing the physical constrains and residual structures could effectively improve the model robustness. Note that although the hidden layers of PI-ResUNet are convolutional layers, the inputs and outputs are still linear layers. Therefore, it is still a continuous PINN that uses AD for differentiation instead of the convolutional differentiation adopted in discrete PINNs (Section 2.3.1). Nasir et al. 333 used PINNs for modeling combustion dynamics in rocket engine chambers, where the 2D governing equations of mass, momentum, energy, and species were considered. With a small amount of data, PINNs were demonstrated to properly model the intricate pressure oscillations and combustion processes.
@@
Figure 10: PINNs for supersonic combustion. (a) PI-ResUNet for reconstructing the combustion flow field in a scramjet combustor 314, where $r$ indicates the density ($\rho$) and $\mu$ includes both the turbulent and laminar viscosity coefficients. (b) Illustration of the flow field reconstruction problem for RDCs 313. Figures are reprinted with permission.
@@
Wang et al. 313 applied PINNs to reconstruct the combustion field in rotating detonation combustors (RDCs) based on partially observed data. As shown in Figure 10(b), a PINN was built to describe the 1D RDC numerical model and estimate the continuous physical fields, taking the space-time coordinates as inputs and physical fields as outputs. In this study, the observations included only ten sets of points with fixed spatial coordinates, representing measured data with fixed sensors. The studied problem was an inverse problem, where the equations and partial true results were known, but the overall solution of the equations was unknown. Results showed that compared to interpolation-based methods, PINN could better capture different distribution patterns for the flow characteristics before and after the detonation wave, which matched the ground truth satisfactorily. The proposed PINN can serve as a surrogate model of the RDC, which has promising engineering value in the design optimization and real-time monitoring of RDCs. After this study, Wang et al. 318, 319 further used the discrete PINNs for SR in RDCs, which will be discussed in Section 5.5.